OpenStudy (anonymous):

Use a matrix to solve the system of equations. x + 5y = -3 -3x + y = 5

hero (hero):

y = (-3-x)/5 y = 5+3x y = y 5(5+3x) = -3 - x 25 + 15x = -3 -x 16x = -28 x = -28/16 = -7/4 y = good luck finding it

hero (hero):

y = 20/4 - 21/4 = -1/4

hero (hero):

GT, stop following me around

OpenStudy (anonymous):

I think using the matrix approach requires using cramer's rule. http://en.wikipedia.org/wiki/Cramer's_rule Look at the section on "Explicit formulas for small systems" that shows how to use it for linear systems with two variables.